# Second- and third-order properties of multidimensional Langevin equations

### v12

@@ 8.4. Dynamics on long time-scales
We now introduce another time-lag $\tau^{\prime}$ and calculate using Eq. (374):
@@
We now investigate the contributions of the various terms in Eq. (384) to $\langle(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$ for $\tau\ll\mathcal{O}(\mu)$ [^18] If we immediately put $\tau\sim\mathcal{O}(\mu)$, then by Eqs. (171) and (172), we will introduce a cubic term in $\langle\widetilde{x}^{i}(\tau)\widetilde{x}^{j}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle-\langle\widetilde{x}^{i}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle\langle\widetilde{x}^{j}(\tau)\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$, which significantly complicates matters.. We do this by multiplying by 1, $\widetilde{\mathbf{x}}(0)$, or $\widetilde{\mathbf{v}}(0)$, taking the expectation and integrating. First, we address the multiplication by 1. The last line of Eq. (384) contributes $\mathcal{O}(\mu^{-1}\tau)$, while the first three lines of Eq. (384) contribute $\mathcal{O}((\mu^{-1}\tau)^{2})$. The remaining terms contribute $\mathcal{O}(\mu^{0})a\mathcal{O}(a,b)\tau$. Thus, only the last line of Eq. (384) survives. Now, we address the multiplication by $\widetilde{x}^{m}(0)$. The last line of Eq. (384) contributes $\mathcal{O}(\mu^{-1}\tau)b$ (to be calculated later), while the first three lines of Eq. (384) contribute $\mathcal{O}((\mu^{-1}\tau)^{2})\mathcal{O}(a,b)$. The fourth line of Eq. (384) contributes:
@@
Note that $\breve{b}_{v}$ computed from Eq. (386) is of the same order as the corresponding quantity computed from the remaining terms in Eq. (384). Now, we again introduce $\tau^{\prime}$ and compute:
@@ 12. Appendix B
(413) $\lim_{E\to\infty}\lim_{\tau\to 0}\int_{-E}^{E}\mathrm{d}x\,p(x)x\frac{\langle x(\tau)-x(0)\mid x(0)=x\rangle}{\tau}\neq\lim_{\tau\to 0}\lim_{E\to\infty}\int_{-E}^{E}\mathrm{d}x\,p(x)x\frac{\langle x(\tau)-x(0)\mid x(0)=x\rangle}{\tau}.$

### v13

@@ 8.4. Dynamics on long time-scales
We now introduce another time-lag $\tau^{\prime}>0$ and calculate using Eq. (375):
@@
We now investigate the contributions of the various terms in Eq. (385) to $\langle(\widetilde{x}^{i}(\tau)-\widetilde{x}^{i}(0))(\widetilde{x}^{j}(\tau)-\widetilde{x}^{j}(0))\mid\widetilde{\mathbf{x}}(0),\widetilde{\mathbf{v}}(0)\rangle$ for $\tau\ll\mathcal{O}(\mu)$ [^18] If we immediately put $\tau\sim\mathcal{O}(\mu)$, then by Eqs. (171) and (172), we will introduce a cubic term, which significantly complicates matters.. We do this by multiplying by 1, $\widetilde{\mathbf{x}}(0)$, or $\widetilde{\mathbf{v}}(0)$, taking the expectation and integrating. First, we address the multiplication by 1. The last line of Eq. (385) contributes $\mathcal{O}(\mu^{-1}\tau)$, while the first three lines of Eq. (385) contribute $\mathcal{O}((\mu^{-1}\tau)^{2})$. The remaining terms contribute $\mathcal{O}(\mu^{0})a\mathcal{O}(a,b)\tau$. Thus, only the last line of Eq. (385) survives. Now, we address the multiplication by $\widetilde{x}^{m}(0)$. The last line of Eq. (385) contributes $\mathcal{O}(\mu^{-1}\tau)b$ (to be calculated later), while the first three lines of Eq. (385) contribute $\mathcal{O}((\mu^{-1}\tau)^{2})\mathcal{O}(a,b)$. The fourth line of Eq. (385) contributes:
@@
Note that $\breve{b}_{v}$ computed from Eq. (387) is of the same order as the corresponding quantity computed from the remaining terms in Eq. (385). Now, we again introduce $\tau^{\prime}>0$ and compute:
@@ 12. Appendix B
(414) $\lim_{E\to\infty}\lim_{\tau\to 0^{+}}\int_{-E}^{E}\mathrm{d}x\,p(x)x\frac{\langle x(\tau)-x(0)\mid x(0)=x\rangle}{\tau}\neq\lim_{\tau\to 0^{+}}\lim_{E\to\infty}\int_{-E}^{E}\mathrm{d}x\,p(x)x\frac{\langle x(\tau)-x(0)\mid x(0)=x\rangle}{\tau}.$